OpenStudy (anonymous):

An air conditioner removes 4.6×10^5 J/min of heat from a house and exhausts 7.8×10^5 J/min to the hot outdoors. How much power does the air conditioner's compressor require?

OpenStudy (anonymous):

So I'm not really sure where to start with this problem. Energy isn't heat so... yeah.

OpenStudy (anonymous):

*pokes through for useful equations that might relate*

OpenStudy (anonymous):

What are you looking for?

OpenStudy (anonymous):

are you looking for efficiency?

OpenStudy (anonymous):

Where to start... oops, question is "How much power does the air conditioner's compressor require?"

OpenStudy (anonymous):

I don't know what that means or how it relates to the "joules of heat" in the air. It seems weird to me.

OpenStudy (anonymous):

What are the units of value above?? Joule? or Calory?

OpenStudy (anonymous):

Um, they're in J/min. :6

OpenStudy (anonymous):

Ah! I didn't mean to close the question... I was just fixing what it said... (I don't use this site much.)

OpenStudy (anonymous):

I fixed it.

OpenStudy (anonymous):

using P = W/t., where P is Power, W is work, and t in time

OpenStudy (anonymous):

W is also Heat, work and so on..,

OpenStudy (anonymous):

Okay, so the work is the heat removed? Does the exhaust cost anything work wise?

OpenStudy (anonymous):

P = ((4.6×10^5 (+ MAYBE 7.8×10^5))*60)/t ~ I'm not sure what t is in this case but I feel like this should result in a rate or something like s^-1/Watts

OpenStudy (anonymous):

i think that's addtion of both

OpenStudy (anonymous):

Multiplying the top by 60 converts the minutes to seconds, seems reasonable.

OpenStudy (anonymous):

work=7.8*10^5

OpenStudy (anonymous):

so...., if @03225186213 is right.., P is W/t = (7.8*10^5/ 60s) Watt

OpenStudy (anonymous):

So W would be the J/min exhausted into the environment by that logic. It seems like the units aren't properly preserved.

OpenStudy (anonymous):

The joules need to be eliminated somehow.

OpenStudy (anonymous):

Can't just drop things... Watts should be s^-1, so the joules need to be eliminated in the formula or by a conversion factor. Assuming power is in Watts.

OpenStudy (anonymous):

s^-1 is frekuensi...., do you remember f = n/t ??

OpenStudy (anonymous):

i mean frequency is s^-1.., sorry

OpenStudy (anonymous):

No... I don't remember. f = n/t Is f frequency? t is definitely time.

OpenStudy (anonymous):

yes f is frequency...,

OpenStudy (anonymous):

AH! You're right! Watts are J/s and Hz are s^-1. (Feels a lot less confused)

OpenStudy (anonymous):

yes another unit of frequency is s^-1 and J/s for Watt

OpenStudy (anonymous):

So I just need to convert the energy taken to make the air cool, that is the exhausts and convert it into Watts. Wow. That's significantly less confusing.

OpenStudy (anonymous):

Which would be... 46,800,000 W? (That doesn't seem right...)

OpenStudy (anonymous):

Hm.... Maybe energy in - energy out?

OpenStudy (anonymous):

Nope... -_- What's t?

OpenStudy (anonymous):

maybe that is eficiency where in/out times 100%

OpenStudy (anonymous):

Efficiency's the next part, but I didn't have any trouble with that part of my homework. It should just be work in/work out.

OpenStudy (anonymous):

Well 1 - that.

OpenStudy (anonymous):

So maybe that means power in this case is just work? Since we're assuming it's perfectly efficient?

OpenStudy (anonymous):

Except... it's not.

OpenStudy (anonymous):

that's work per sekon, so that's power..., It's just that in the view of work

OpenStudy (anonymous):

I wanna say 4.6×10^5 J/min + 7.8×10^5 J/min And if you want it in watts just convert from J/min to j/ s Also I think capital J means thousands where as j is just a joule

OpenStudy (anonymous):

Sekon isn't a unit I'm familiar with...

OpenStudy (anonymous):

yes i think so @timo86m ...,

OpenStudy (anonymous):

sory i mean seconds

OpenStudy (anonymous):

Ah, that explains why I was so confused. ^_^

OpenStudy (anonymous):

20667 W doesn't work....

OpenStudy (anonymous):

may I see your question paper?

OpenStudy (anonymous):

It's not on paper. :P Question posted is all there is. It's online.

OpenStudy (anonymous):

So the top is all the information given.

OpenStudy (anonymous):

So what's your choice?

OpenStudy (anonymous):

What do you mean choice?

OpenStudy (anonymous):

The choice is to keep working on it or not... I'm thinking about it and experimenting with my given values...

OpenStudy (anonymous):

I guess I'll just work on it offline though. I don't seem to be making much progress online. I'm sorry I bothered you guys.

OpenStudy (anonymous):

it's your decision ^^

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

hey you no bother at all :)

OpenStudy (anonymous):

Ahahaha! I figured it out! It was work out - work in! =D

OpenStudy (anonymous):

that is odd

OpenStudy (anonymous):

Nah, it makes sense. :) It's doing work on the air so the air's energy doesn't count towards the energy used by it being plugged in.

OpenStudy (anonymous):

i got 20666.66667 watts or 20.666 Kilo watts sounds reasonable

OpenStudy (anonymous):

Yup, that's what you get if you add them. The energy in the air makes it so the electricity doesn't have to do as much work.